OpenStudy (anonymous):

5. For this question, you will need a parent/guardian or a friend. Have this individual grab a handful of coins making sure there are only two types of coins in the group (i.e., nickels and dimes, quarters and pennies, pennies and dimes, etc). Your parent/guardian or friend should tell you the type of coins they’ve chosen, how many coins they have and the dollar amount of the group. From this information, you will set up two sets of equations and determine how many of each coin they have in their hand.

OpenStudy (anonymous):

q + p = 15 -q -q p = -q + 15 .25q +. 01( -q + 15)=.39 .25q - .01q + .15 =.39 .24q + .15 = .39 - .15 -.15 ------------- .24q = .24 ----- ---- .24 .24 q = 1 p = -q + 15 p = -1 + 15 p = 14